OpenStudy (anonymous):

given the function p decribed by p(x)=x^3-x, find p(-8)

OpenStudy (nory):

p(x) = x^3 - x p(-8) = (-8) ^3 - (-8) Now evaluate that. What do you get?

OpenStudy (imtiaz7):

p(x)=x^3-x, find p(-8) by putting the value of x which is -8: p(-8)= (-8)^3-(-8) p(-8)= -512+8 p(-8)= 504

OpenStudy (imtiaz7):

there is a -504

OpenStudy (anonymous):

Thank you both. Imtiaz7 especially, that was what I was looking for. Have a good one

OpenStudy (imtiaz7):

no problem mygumshurt.

OpenStudy (nory):

Got any other problems?

OpenStudy (anonymous):

That's it. Thanks again. Be well